OpenStudy (anonymous):

two pyramids below are similar.What is the length of the smaller pyramid? 5;2 3;x

OpenStudy (saifoo.khan):

5x = 6 x = 6/5

OpenStudy (anonymous):

use the cross product method (like is shown above) in future problems

OpenStudy (anonymous):

like how

OpenStudy (saifoo.khan):

5 into x, 3 into 2

OpenStudy (anonymous):

so will you check my answer?

OpenStudy (saifoo.khan):

Sure.

OpenStudy (anonymous):

5x = 3(2) 5x = 6 x = 6/5

OpenStudy (anonymous):

is it 10/3

OpenStudy (anonymous):

if its the same problem... false